Forums Calendar Gradebook sages Nabandyan 17878> Assessment Du Let f(z)= (5z- 6z) (6+ ) Find f(z). f(z)= Preview TIP Enter
Answer #1

If(n)= (54-6m3) (6+5m) fo=(58-628)6+7)+ [ 64 707 (562) ! f(m) = (5-1872) (6+54) + (8 + 2/1 m) (50 - 643) f(x) = 30 +550-10892